how many sides does a regular polygon have if the sum of its interior angle measures is 3240?
Answer by unlockmath(1688)

Hello,
There is a formula to follow which is:
(n-2)180=total interior angle degrees. So we can plug in the known value:
(n-2)180=3240 Rewritten as:
180n-360=3240 Add 360 to both sides and divide by 180 to get:
n=20
There we go, 20 sides.
Make sense?
RJ
Check out a new book at:
www.math-unlock.com